# APVR: Hour-Level Long Video Understanding with Adaptive Pivot Visual Information Retrieval

### v1

@@ Abstract
Current video-based multimodal large language models struggle with hour-level video understanding due to computational constraints and inefficient information extraction from extensive temporal sequences. We propose APVR (Adaptive Pivot Visual information Retrieval), a training-free framework that addresses the memory wall limitation through hierarchical visual information retrieval. APVR operates via two complementary components: Pivot Frame Retrieval employs semantic expansion and multi-modal confidence scoring to identify semantically relevant video frames, while Pivot Token Retrieval performs query-aware attention-driven token selection within the pivot frames. This dual granularity approach enables processing of hour-long videos while maintaining semantic fidelity. Experimental validation on LongVideoBench and VideoMME demonstrates significant performance improvements, establishing state-of-the-art results for not only training-free but also training-based approaches while providing plug-and-play integration capability with existing MLLM architectures.
@@ 1 Introduction
The proliferation of long-form video content necessitates robust understanding capabilities that extend beyond the current limitations of video-based multimodal large language models. While exist models (Lin et al. 2023; Zhang et al. 2025) demonstrate proficiency on short video sequences, they encounter fundamental scalability challenges when processing hour-level content, where the quadratic complexity of attention mechanisms creates prohibitive computational demands and semantic information becomes diluted across extensive temporal spans.
@@
The core challenge stems from the inability of current Video MLLMs to effectively model and extract critical information from long video sequences. In hour-level videos, semantically relevant content typically represents a sparse subset of the total temporal sequence, yet existing approaches process entire sequences by uniform sampling, leading to computational inefficiency and degraded performance due to attention dispersion across irrelevant segments.
Contemporary approaches to long video understanding bifurcate into training-based and training-free paradigms, each presenting distinct limitations that constrain practical deployment. Training-based methods employ strategies such as multimodal sequence parallelism(Fu et al. 2025; Liu et al. 2024d) or compression techniques (Shu et al. 2024; Liu et al. 2025a) during the training phase to accommodate extended sequences. However, these approaches suffer from several critical drawbacks that limit their applicability. First, they require substantial training datasets and computational resources, making them prohibitively expensive for many new scenes. Second, the long-video training process often compromises the model’s foundational capabilities on shorter sequences, necessitating replay training to maintain performance across diverse sequence lengths. Third, these methods exhibit poor adaptability when confronted with novel task distributions that differ significantly from their training domains, requiring extensive retraining or fine-tuning procedures.
Training-free methods present an alternative paradigm focused on keyframe retrieval and selective processing strategies (Tang et al. 2025a; Cheng et al. 2025a; Luo et al. 2025). While these approaches offer computational advantages and deployment flexibility, they face fundamental limitations in their current implementations. Existing keyframe retrieval methods typically select only a limited number of video frames, which proves insufficient for tasks requiring comprehensive temporal reasoning and long-range logical understanding. Furthermore, when researchers attempt to increase the number of selected frames to improve coverage, these methods quickly encounter out-of-memory limitations, effectively hitting the memory wall that prevents processing of truly long sequences. The trade-off between sequence coverage and computational feasibility has prevented existing training-free methods from achieving the temporal comprehensiveness necessary for complex video understanding tasks.
Our proposed Adaptive Pivot Visual information Retrieval (APVR) framework addresses these limitations through a novel training-free approach that combines strategic keyframe retrieval with adaptive token selection and compression mechanisms. Our approach is motivated by the observation that effective long video understanding requires strategic information retrieval at multiple granularities. Rather than processing entire video sequences uniformly, APVR employs a two-stage retrieval mechanism that operates at both frame and token levels to identify and extract pivot visual information that is most relevant to the input query. This dual granularity strategy enables the system to breakthrough the memory wall while maintaining semantic accuracy across long video understanding tasks. By focusing computational resources on the most semantically relevant frames and subsequently performing intelligent token compression within those frames, APVR achieves comprehensive coverage of hour-long videos without encountering memory limitations.
@@
The core innovation of APVR lies in its adaptive retrieval strategy that combines semantic information expansion with multi-modal confidence scoring. At the frame level, our Pivot Frame Retrieval (PFR) component expands the original query into four types of semantic information: objects, descriptions, relations, and semantics. This expansion enables more comprehensive frame scoring using complementary visual models, specifically CLIP for semantic similarity and Grounding-DINO for object detection and spatial reasoning. The framework incorporates temporal diffusion mechanisms to maintain temporal coherence and employs adaptive resampling strategies to progressively refine frame selection across multiple iterations. At the token level, our Pivot Token Retrieval (PTR) component extends the concept of semantic importance to fine-grained visual representations. This component leverages query-aware multi-layer attention scoring to identify the most relevant visual tokens within selected frames, employing dynamic chunk-wise selection and head-wise soft voting mechanisms to maintain both computational efficiency and semantic accuracy.
@@
$\bullet$ We develop a mechanism that leverages multi-modal confidence scoring and query-aware attention scoring to preserve temporal structure and semantic details, ensuring accurate understanding of complex video narratives while maintaining computational efficiency.
$\bullet$ We demonstrate that intelligent multi-granularity retrieval provides a sustainable alternative to parameter scaling for long video understanding. Our training-free design ensures plug-and-play integration with existing MLLM architectures while preserving foundational capabilities.
@@ 2 Related Work
Long Video Understanding  Recently advances in long video understanding can be broadly categorized into training-based and training-free paradigms. In the training-based paradigm, MLLMs are directly trained using large-scale long video datasets. The whole training process is usually divided into several stages to stably improve different components of MLLM, as well as to improve the ability to model different types of input visual data (Li et al. 2024b; Zhang et al. 2025). In order to process extra-long videos, current methods explore two different strategies: sequence parallelism (Chen et al. 2024; Zhang et al. 2024b; Shen et al. 2025) or visual feature compression (Song et al. 2024; Shen et al. 2024; Liu et al. 2025a). For instance, LongVILA (Chen et al. 2024) adopts a five-stage training pipeline with a novel multi-modal sequence parallelism to efficiently understand long videos. Video-XL-Pro (Liu et al. 2025a) builds a learnable module to generate compact and comprehensive video tokens for hour-long video understanding. In the training-free paradigm, long video understanding task is organized as an agent-like process (Liu et al. 2025b; Wang et al. 2024c), where MLLMs serve as reasoning engine within the system. These training-free methods mainly focus on retrieving the most important information in video frames (Tang et al. 2025b; Park et al. 2024; Ye et al. 2025; Guo et al. 2025; Cheng et al. 2025b) or visual tokens (Wang et al. 2025a; Luo et al. 2025). For example, AKS (Tang et al. 2025b) adopts a plug-and-play module for keyframe selection and information pre-filtering for video-based MLLMs. QuoTA (Luo et al. 2025) adopts an ante-hoc training-free modular for query-aware visual token assignment and important information retrieval. Despite these advancements, few methods simultaneously explore frame and token optimization, which is addressed in this work through adaptive pivot information retrieval.
@@ 3 Method
Figure 1: The overall framework of our proposed APVR. We integrate two plug-and-play components, Pivot Frame Retrieval and Pivot Token Retrieval, into MLLMs to improve the performance of long video understanding. APVR is training-free and provides a alternative to parameter scaling. With frame-level and token-level adaptive selection, it can accurately undertand complex videos with computational efficiency.
@@ 3.1 Overview
The framework of the proposed APVR is illustrated in Fig. 1, which is a training-free system. Given an input video =$\mathcal{F}=\{f_{t}\}_{t=1}^{N}$ with $N$ frames and query $Q$, the system aims to retrieve the most important information to feed into MLLM for the final response. To achieve this, we build two main components in APVR: Pivot Frame Retrieval (PFR) and Pivot Token Retrieval (PTF). In PFR, we design an efficient strategy to search pivot frames $\mathcal{F}_{pivot}=\{f_{t}\}_{t=1}^{K}$ with expanded queries and basic visual models. In PTR, we explore an attention-driven token selection to retrieve pivot tokens $\mathcal{T}_{pivot}$ that are most relevant with expanded queries. Finally, any type of MLLM can be integrated into our plug-and-play APVR to generate accurate response to the input video and query.
@@ 3.2 Pivot Frame Retrieval
We build an efficient algorithm to retrieve the pivot frames in an iterative manner. To be specific, we first expand the original query into four types of semantic information using LLM. The expanded information is then utilized for frame scoring using two basic visual models to obtain different types of confidence. Finally, all confidence is properly integrated to sort the video frames and select the top-K frames for the subsequent module.
@@
To filter out the most relevant video frame only based on the brief query is challenging. Thus, we expand the given query $Q$ to four types of semantic information as follows:
@@
Objects: The key objects $O_{k}$ or cue objects $O_{c}$ that are detectable by visual models such as Grounding-DINO (e.g., ”bold person”,”phone”,”blue background”).
@@
Descriptions: The entity or hypernym concepts of objects based on knowledge graph augmentation, denoted as $D$. (e.g. ”pearl headband: decorative accessory worn on head; jewelry”)
@@
Relations: The triplet relationship $R\subseteq O\times\mathcal{R}\times O$ to enhance the confidence of frames with logical links between objects. In each triplet $(o_{i},r,o_{j})$, the type of relation $r\in\mathcal{R}$ is exactly one of the following: (1) spatial which means that $o_{j}$ and $o_{j}$ appear in the same frame such as ”a bold man holding a red phone”. (2) time which means that $o_{j}$ and $o_{j}$ appear in different frames in orders such as ”A woman walks through, after a cat follows.” (3) attribute which means that $o_{i}$ appears with an attribute describable by $o_{j}$ such as ”A woman in a white skirt”. (4) causal which means that $o_{j}$ and $o_{j}$ appear in orders with a cause-effect manner such as ”A woman opens the door, the cat is scared and runs out.”
@@
Semantics: The semantics information of query and options based on knowledge graph, denoted as $M$ (e.g., leash often appears with dog).
Multi-modal Confidence Scoring
@@
Given a video $\mathcal{F}$ consisting of $N$ frames, and a textual query $Q$ with expanded semantic information $\{O_{k},O_{c},D,R,M\}$, instead of exhaustively processing all frames, we first uniformly sample few frames with stride $\nabla$ in the first turn scoring and then adaptively resample frames for the subsequent iteration based on the online updated frame confidence.
@@
Our algorithm exploits two basic visual models, i.e., Grounding-DINO and CLIP, to comprehensively score the confidence of video frames and iteratively optimize the overall confidence as follows:
@@
CLIP-based Similarity Scoring: We leverage the CLIP model to compute cross-modal similarity between text and visual features. For each of the sampled frames $\{f_{t}\}_{t=1}^{N_{s}}$, we encode it using the CLIP image encoder $\phi_{I}(\cdot)$:
$v_{t}=\frac{\phi_{I}(f_{t})}{||\phi_{I}(f_{t})||_{2}}.$ (1)
The query as well as the expanded description and semantics are aggregated and encoded using the CLIP text encoder $\phi_{T}(\cdot)$:
$t_{agg}=\sum_{e}\frac{\phi_{T}(e)}{||\phi_{T}(e)||_{2}},e\in\{Q,D,M\}.$ (2)
The CLIP similarity score for frame $f_{t}$ is then calculated as:
$s_{t}^{CLIP}=softmax(\tau\cdot v_{t}\cdot t_{agg})$ (3)
, where $\tau$ is a temperature parameter set to $100$ to sharpen the score distribution.
Grounding-DINO Object Detection: To enhance the retrieval precision with key/cue objects along with their relations, we first employ Grounding-DINO to identify extracted objects:
$\mathcal{B}_{t},\mathcal{L}_{t}=\Psi_{GD}(f_{t},O_{k}\cup O_{c})$ (4)
, where $\Psi(\cdot)$ is the Grounding-DINO model that returns bounding boxes $\mathcal{B}_{t}$ and corresponding confidence logits $\mathcal{L}_{t}$ for frame $f_{t}$. The score of detection is then computed as:
$s_{t}^{o}=softmax(max(\mathcal{L}_{t})).$ (5)
To further model the logic relations among objects in the same or different frames to enhance the detection results, we update the score by adding corresponding frame score if there are satisfied relation triplets in this frame as follows:
$s_{t}^{GD}=s_{t}^{o}+\sum_{r}(\omega_{r}\cdot s_{r})$ (6)
, where $r$ is exactly one of spatial/time/attribute/causal and $\omega_{r}$ is the corresponding weight to control the contribution of each type of relation.
@@
Temporal Diffusion and Adaptive Resampling
To account for the temporal continuity of video content and improve frame resampling, we introduce a temporal score diffusion mechanism that propagates confidence scores across neighboring frames. Specifically, for a scored frame $t$, we update the score distribution in a window:
$\mathcal{S}_{i}=max(\mathcal{S}_{i},\frac{\mathcal{S}_{t}}{1+|i-t|}),i\in[t-w,t+w]$ (8)
This diffusion process ensures that frames adjacent to high-confidence frames receive boosted scores proportional to their proximity, which helps capture temporally coherent segments relevant to the query. The score distribution of the whole video $\mathcal{S}$ is then updated. In addition, the frames scored in iteration $p$ are marked as visited $\{\epsilon_{t}=1\mid f_{t}\in\mathcal{F}_{vis}^{p}\}$. Based on $\mathcal{S}$, we further design an adaptive and hybrid candidate sampling strategy for subsequent iterations that combines two types of set:
High-confidence Set: We select frames with higher scores from all unvisited frame set $\mathcal{C}_{t}=\{f_{t}\mid\epsilon_{t}=0\}^{N_{c}}_{t=1}$ based on the distribution as:
$\mathcal{H}_{t}=topk(\mathcal{S}\odot\mathcal{C}_{t}),k=\frac{N}{2\nabla}$ (9)
Uncertainty Set: We identify the regions with high Shannon entropy in the score distribution as:
$\mathcal{E}_{\gamma}(s_{i})=-\sum_{j=i-\gamma}^{i+\gamma}e_{j}log(e_{j}),e_{j}=\frac{s_{j}}{\sum_{k=j-\gamma}^{j+\gamma}s_{k}}$ (10)
, then the sampled frame is determined by:
$\mathcal{U}_{t}=\{f_{t}\mid\mathcal{E}_{\gamma}(s_{t})>\mu+0.5\sigma\}$ (11)
, where $\mu$ and $\sigma$ are the mean and standard deviation of entropy values across candidate set.
The final union candidate set is $\mathcal{L}_{t}=\mathcal{H}_{t}\cup\mathcal{U}_{t}$, which will be used for adaptive sampling. To ensure progressively finer sampling, we decrease the stride in each iteration by $\nabla^{p}=max(1,\frac{\nabla}{p+1}$). The selected frames $L_{selected}$ for next iteration are sample in the way as:
$\mathcal{L}_{selected}=L_{*}\cup L_{rand},$ (12)
$L_{*}=Multinomia(S_{\mathcal{L}_{t}},(1-\alpha)N_{c}),$ (13)
$L_{rand}=U(Perm(\alpha N_{c}))$ (14)
, where $\alpha$ balances the score-based importance sampling $L_{*}$ and the diverse random sampling $L_{rand}$ to capture the full semantic content of the video and also avoid falling into local optimum.
After $P$ iterations, the $K$ pivot frames are retrieved by $\mathcal{F}_{pivot}=topk(\mathcal{S}\odot\mathcal{F})=\{f_{t}\}_{t=1}^{K}$ and fed into the next module.
@@
ViLaMP(Cheng et al. 2025a) 7B training-based 57.8 - - - 67.5
@@
Figure 2: Qualitative Comparison of APVR with the baseline MLLM. The expanded semantic information is significant complementary for pivot frame retrieval. The number and score of the pivot frame is drawn in the left-top with green and red text, respectively. Yellow stars indicate the key frame for answering the question. Green answer represents correct while red one represents wrong.

### v2

@@ Abstract
Current multimodal large language models (MLLMs) struggle with hour-level video understanding, facing significant challenges not only in modeling the substantial information volume of long videos but also in overcoming the memory wall and resource constraints during both training and inference. Although recent training-free approaches have alleviated resource demands by compressing visual features, their reliance on incomplete visual information limits the performance potential. To address these limitations, we propose Adaptive Pivot Visual information Retrieval (APVR), a training-free framework that hierarchically retrieves and retains sufficient and important visual information. It breakthroughs the memory wall limitation via two complementary components: Pivot Frame Retrieval employs query expansion and iterative spatio-semantic confidence scoring to identify relevant video frames, and Pivot Token Retrieval performs query-aware attention-driven token selection within up to 1024 pivot frames. This dual granularity approach enables the processing of hour-long videos while maintaining semantic fidelity. Experimental validations demonstrate significant performance improvements, achieving 64.9% on LongVideoBench and 68.4% on VideoMME, which are state-of-the-art results for both training-free and training-based approaches. Meanwhile, our method provides plug-and-play integration capability with existing MLLM architectures.
@@ 1 Introduction
The proliferation of long-form video content necessitates robust understanding capabilities that extend beyond the current limitations of multimodal large language models (MLLMs). While existing video-based models (Lin et al. 2023; Zhang et al. 2025) demonstrate proficiency on short video sequences, their scalability for hour-level video remains challenging. Specifically, they encounter not only the memory wall but also the inherent dilution of semantic information across extensive temporal spans.
@@
The core challenge stems from the inability of current Video MLLMs to effectively model and extract critical information from long video sequences. In hour-level videos, semantically relevant contents are typically sparse across both temporal and spatial dimensions. However, existing approaches process entire sequences by uniform sampling, leading to computational inefficiency and degraded performance due to attention dispersion across irrelevant segments. As illustrated in Fig. 1(a), this strategy samples frames that are not pivot, leading to wrong answer generated by MLLMs.
Figure 1: Given a long video, uniform sampling (a) and insufficient frame retrieval (b) yield incorrect answer, while naively increasing pivot frames (c) raises OOM. Our APVR (d) explores joint frame-token co-retrieval, concentrating computational resources on the most relevant information.
Recently, some approaches (Ataallah et al. 2024; Tang et al. 2025a; Guo et al. 2025; Ye et al. 2025; Cheng et al. 2025; Luo et al. 2025) have attempted to retrieve query-relevant frames from long videos, simplifying long video understanding task as key-frame selection. Consequently, MLLMs process only few and sparse key frames instead of continuous video frames, neglecting the temporal-semantic relationships and potentially failing in scenarios requiring comprehensive temporal reasoning and long-range logical understanding, as shown in Fig. 1(b).
A potential solution to the above limitation is to retrieve sufficient pivot frames and improve coverage as shown in Fig. 1(c). However, such approaches encounter the memory wall when processing these frames with MLLMs. While recent training-based methods employ strategies such as multimodal sequence parallelism(Fu et al. 2025; Liu et al. 2024d) or feature compression (Shu et al. 2024; Liu et al. 2025a) during the training phase to accommodate extended video frames, they demand resource-intensive multi-stage training based on existing MLLMs. Thus, they lack adaptability to rapidly evolving MLLM architectures due to high retraining costs.
The inherent trade-off between temporal coverage and computational feasibility has hindered existing methods from achieving comprehensive video understanding. To address this core challenge, we propose Adaptive Pivot Visual information Retrieval (APVR) framework, a training-free approach that integrates spatio-semantic frame retrieval with token retrieval and compression, as shown in Fig. 1(d). This two-stage retrieval mechanism enables the system to concentrate computational resources on the pivot visual information that is most relevant to the input query, thus efficiently processing hour-long videos without encountering memory limitations.
@@
Specifically, at the frame level, our Pivot Frame Retrieval (PFR) component expands the original query into four types of semantic information: objects, descriptions, relations, and semantics. This expansion enables more comprehensive frame scoring from spatial and semantic perspectives using complementary visual models: CLIP for semantic similarity and Grounding-DINO for object detection and spatial reasoning. The framework incorporates temporal diffusion mechanisms to maintain temporal coherence and employs adaptive resampling strategies to progressively refine frame selection across multiple iterations. At the token level, our Pivot Token Retrieval (PTR) component extends the concept of semantic importance to fine-grained visual representations. This component leverages query-aware multi-layer attention scoring to identify the most relevant visual tokens within selected frames, employing dynamic chunk-wise selection and head-wise soft voting mechanisms to maintain both computational efficiency and semantic accuracy.
@@
$\bullet$ We develop a mechanism that leverages spatio-semantic confidence scoring and query-aware attention scoring to preserve temporal structure and semantic details, ensuring accurate understanding of complex video narratives while maintaining computational efficiency.
$\bullet$ We demonstrate that intelligent dual-granularity retrieval provides a sustainable alternative to parameter scaling for long video understanding. Our training-free design ensures plug-and-play integration with existing MLLM architectures while preserving foundational capabilities.
@@ 2 Related Work
Long Video Understanding  Recently advances in long video understanding can be broadly categorized into training-based and training-free paradigms. In the training-based paradigm, MLLMs are directly trained using large-scale long video datasets. The whole training process is usually divided into several stages to stably improve different components of MLLM, as well as to improve the ability to model different types of input visual data (Li et al. 2024b; Zhang et al. 2025). In order to process extra-long videos, current methods explore two different strategies: sequence parallelism (Chen et al. 2024; Zhang et al. 2024b; Shen et al. 2025) or visual feature compression (Song et al. 2024; Shen et al. 2024; Liu et al. 2025a). For instance, LongVILA (Chen et al. 2024) adopts a five-stage training pipeline with a novel multi-modal sequence parallelism to efficiently understand long videos. Video-XL-Pro (Liu et al. 2025a) builds a learnable module to generate compact and comprehensive video tokens for hour-long video understanding. In the training-free paradigm, long video understanding task is organized as an agent-like process (Liu et al. 2025b; Wang et al. 2024c), where MLLMs serve as reasoning engine within the system. These training-free methods mainly focus on retrieving the most important information in video frames (Tang et al. 2025b; Park et al. 2024; Ye et al. 2025; Guo et al. 2025; Cheng et al. 2025) or visual tokens (Wang et al. 2025a; Luo et al. 2025). For example, AKS (Tang et al. 2025b) adopts a plug-and-play module for keyframe selection and information pre-filtering for video-based MLLMs. QuoTA (Luo et al. 2025) adopts an ante-hoc training-free modular for query-aware visual token assignment and important information retrieval. Despite these advancements, few methods simultaneously explore frame and token optimization, which is addressed in this work through adaptive pivot information retrieval.
@@ 3 Method
Figure 2: The overall framework of our proposed APVR. We integrate two plug-and-play components, Pivot Frame Retrieval and Pivot Token Retrieval, into MLLMs to improve the performance of long video understanding. APVR is training-free and provides a alternative to parameter scaling. With frame-level and token-level adaptive selection, it can accurately understand complex videos with computational efficiency.
@@ 3.1 Overview
The framework of the proposed APVR is illustrated in Fig. 2, which is a training-free system. Given an input video =$\mathcal{F}=\{f_{t}\}_{t=1}^{N}$ with $N$ frames sampled under specific $fps$ and query $Q$, the system aims to retrieve the most important visual information for accurate response. To achieve this, we build two main components in APVR: Pivot Frame Retrieval (PFR) and Pivot Token Retrieval (PTF). In PFR, we search pivot frames $\mathcal{F}_{pivot}$ with high confidence scores using expanded queries and basic visual models. In PTR, we explore an attention-driven token selection in MLLMs to retrieve pivot tokens $\mathcal{T}_{pivot}$ that are most relevant with expanded queries.
@@ 3.2 Pivot Frame Retrieval
The progress of PFR is developed in an iterative and efficient manner, as elaborated in Alg. 1.
@@
First, We expand the given query $Q$ to four types of semantic information as follows:
@@
Objects: The key objects $Obj_{k}$ or cue objects $Obj_{c}$ that are detectable by visual models such as Grounding-DINO (e.g., “bold person”,“phone”,“blue background”’). All objects are aggregated as $Obj=\{Obj_{k},Obj_{c}\}$.
@@
Descriptions: The entities or hypernym concepts of objects based on knowledge graph augmentation, denoted as $Des$. (e.g. “pearl headband: decorative accessory worn on head; jewelry”)
@@
Relations: The triplet relationship $Rel\subseteq Obj\times\mathcal{R}\times Obj$ to enhance the confidence of frames with logical links between objects. In each triplet $(o_{i},r,o_{j})$, the type of relation $r\in\mathcal{R}$ is exactly one of the following: (1) spatial which means that $o_{j}$ and $o_{j}$ appear in the same frame, such as “a bold man holding a red phone”. (2) time which means that $o_{j}$ and $o_{j}$ appear in different frames in orders such as “A woman walks through, after a cat follows.” (3) attribute which means that $o_{i}$ appears with an attribute describable by $o_{j}$ such as “A woman in a white skirt”. (4) causal which means that $o_{j}$ and $o_{j}$ appear in orders with a cause-effect manner such as “A woman opens the door, the cat is scared and runs out.”
@@
Semantics: The semantics information of query and options based on knowledge graph, denoted as $Sem$ (e.g., ‘leash often appears with dog”).
The prompt template for LLM-based query expansion is provided in Appendix.
Algorithm 1 Pivot Frame Retrieval (PFR)
1:video frames $\mathcal{F}$, query $Q$, iterations $P$, stride $\nabla$
2:Pivot frames $\mathcal{F}_{pivot}=\{f_{t}\}_{t=1}^{K}$
3:Initialize:
4:$\mathcal{S}\leftarrow\mathbf{0}^{N}$triangleright Confidence scores
5:$\mathcal{V}\leftarrow\emptyset$triangleright Visited frames set
6:$\{Obj,Des,Rel,Sem\}\leftarrow\operatorname{LLM}(Q)$triangleright Query Expansion
7:for $p=1$ to $P$ do
8:     $\nabla_{p}\leftarrow\max\left(1,\frac{\nabla}{p}\right)$
9:     $\mathcal{F}_{samp}^{p}\leftarrow\operatorname{AdaptiveResample}(\mathcal{F},\mathcal{S},\mathcal{V},\nabla_{p})$
10:     for $f_{t}\in\mathcal{F}_{samp}^{p}$ do
11:         $s^{CLIP}_{t}\leftarrow\Pi_{CLIP}(f_{t},Q,Des,Sem)$
12:         $s^{GD}_{t}\leftarrow\Pi_{GD}(f_{t},Obj,Rel)$
13:         $\mathcal{S}_{t}\leftarrow(1-\lambda)s^{CLIP}_{t}+\lambda s^{GD}_{t}$
14:         $\mathcal{S}_{[t-w,t+w]}\leftarrow\max\left(\mathcal{S}_{[t-w,t+w]},\frac{\mathcal{S}_{t}}{1+|i-t|}\right)$
15:     end for
16:     $\mathcal{V}\leftarrow\mathcal{V}\cup\mathcal{F}_{samp}^{p}$
17:end for
18:return $\mathcal{F}_{pivot}\leftarrow\operatorname{topk}(\mathcal{S}\odot\mathcal{F},K)$
Spatio-Semantic Confidence Scoring
@@
Given a video $\mathcal{F}$, and query $Q$ with expanded semantic information $\{Obj,Des,Rel,Sem\}$, instead of exhaustively processing all $N$ frames, we uniformly sample few frames with initial stride $\nabla$ in the first iteration and adaptively resample frames for the subsequent iteration based on the online updated frame confidence. To ensure progressively finer sampling, we gradually decrease the stride by $\nabla^{p}=\mathrm{max}(1,\frac{\nabla}{p}$). The frames sampled in iteration $p$ are denoted as $\mathcal{F}_{samp}^{p}$.
@@
Our algorithm exploits two basic visual models, i.e., Grounding-DINO and CLIP, to comprehensively score $\mathcal{F}_{samp}^{p}$ as follows:
@@
CLIP-based Similarity Scoring: We leverage the CLIP model to compute cross-modal similarity between text and visual features as the semantic confidence score based on expanded query and the visual feature by $s_{t}^{CLIP}=\Pi_{CLIP}(f_{t},Q,Des,Sem)$. The details of the function $\Pi_{CLIP}$ are provided in Appendix.
Grounding-DINO Object Detection: We leverage the Grounding-DINO model to compute the spatial confidence score based on the query-aware objects long with their relations by $s_{t}^{GD}=\Pi_{GD}(f_{t},Obj,Rel)$. The details of the function $\Pi_{GD}$ are provided in Appendix.
@@
The frames scored in the current iteration are put into the visited frame set $\mathcal{V}$.
Temporal Diffusion: To account for the temporal continuity of video content and improve frame resampling, we introduce a temporal score diffusion mechanism that propagates confidence scores across neighboring frames as:
$\mathcal{S}_{i}=\mathrm{max}(\mathcal{S}_{i},\frac{\mathcal{S}_{t}}{1+|i-t|}),i\in[t-w,t+w]$ (2)
This diffusion process helps capture temporally coherent segments relevant to the query.
Adaptive Resampling
Based on the updated $\mathcal{S}$, we further design an adaptive and hybrid candidate sampling strategy for subsequent iterations that combines two types of set:
High-confidence Set: We select frames with higher scores from all unvisited frame set $\mathcal{U}=\{f_{t}\mid f_{t}\notin\mathcal{V}\}$ based on the distribution as:
$\mathcal{H}_{s}=\mathrm{TopK}(\mathcal{S}\odot\mathcal{U},\frac{N}{2\nabla_{p}})$ (3)
Uncertainty Set: We calculate Shannon entropy based on the score distribution as:
$\mathcal{E}_{\gamma}(s_{i})=-\sum_{j=i-\gamma}^{i+\gamma}e_{j}\cdot\text{log}(e_{j}),e_{j}=\frac{s_{j}}{\sum_{k=j-\gamma}^{j+\gamma}s_{k}}$ (4)
, then the frames with high entropy are identified by:
$\mathcal{H}_{e}=\{f_{t}\mid\mathcal{E}_{\gamma}(s_{t})>\mu+0.5\sigma\}$ (5)
, where $\mu$ and $\sigma$ are the mean and standard deviation of the entropy in $\mathcal{C}$.
The union candidate set for this iteration sampling is $\mathcal{C}=\mathcal{H}_{s}\cup\mathcal{H}_{e}$. Assume that there are $N_{\mathcal{C}}$ frames in $\mathcal{C}$. The frames $\mathcal{F}_{samp}^{p}$ are sampled in the way as:
$\mathcal{F}_{samp}^{p}=\mathcal{C}_{\text{multi}}\cup\mathcal{C}_{\text{rand}},$ (6)
$\mathcal{C}_{\text{multi}}\sim\mathrm{Multinomial}(S\odot\mathcal{C},(1-\alpha)N_{\mathcal{C}}),$ (7)
$\mathcal{C}_{\text{rand}}\sim\mathrm{Unif}(\mathrm{Perm}(S\odot\mathcal{C},\alpha N_{\mathcal{C}}))$ (8)
, where $\alpha$ balances the multinomial score-based importance sampling $\mathcal{C}_{\text{multi}}$ and the diverse random sampling $\mathcal{C}_{\text{rand}}$ to capture the full semantic content of the video and also to avoid falling into local optimal.
After $P$ iterations, the $K$ pivot frames are retrieved by $\mathcal{F}_{pivot}=\mathrm{TopK}(\mathcal{S}\odot\mathcal{F},K)$ and fed into the next module.
@@
ViLaMP(Cheng et al. 2025) 7B training-based 57.8 - - - 67.5
@@
Figure 3: Qualitative Comparison of APVR with the baseline MLLM. The expanded query is significant complementary for pivot information retrieval. The number and score of the pivot frame is drawn in the left-top with green and red text, respectively. Yellow stars indicate the key frame for correct response. Green answer represents correct while red one represents wrong.